# Evaluating the Efficacy of Prompt-Engineered Large Multimodal Models Versus Fine-Tuned Vision Transformers in Image-Based Security Applications

### v1

@@ 5.1.3 Evaluation Metrics
The performance of the Gemini-pro-vision and fine-tuned ViT models was assessed using standard classification metrics, including accuracy, precision, recall, F1-score, and confusion matrix. These metrics provided a comprehensive understanding of each model’s efficacy in correctly identifying triggered images and classifying malware types and families.
  • 1.
@@
  • 3.
    Recall (or sensitivity) measures the proportion of actual positives correctly identified.
  • 4.
@@ 5.2 Task 1: Trigger Detection
5.2.1 Prompt engineering with Gemini-pro
We implement the methodology outlined earlier using the Gemini-pro model on the MNIST dataset, experimenting with different input prompts to assess their impact on model performance. Our investigation focuses on three distinct prompt templates for detecting triggers in images, as illustrated in Figure 4.
  • 1.
    Prompt 1: The baseline prompt asks the LMM to classify images as ’Triggered’ or ’Unaltered’ without further details. This prompt led to all images being classified as ’Unaltered,’ resulting in a 50% accuracy rate. This outcome demonstrates a significant limitation in the LMM’s ability to identify triggers without specific guidance. The model’s default behavior to opt for the safer ’Unaltered’ classification suggests a conservative approach in ambiguous scenarios, highlighting the need for more descriptive and directive prompts in cybersecurity applications.
  • 2.
    Prompt 2: By enhancing the prompt to define a trigger as a small white square, the model’s performance improved with an accuracy of 63.00% and F1-Score of 45.10%. However, the low recall of 30.40% indicates a tendency to miss many true ’Triggered’ cases. This suggests that while providing a description of the trigger helps, it is insufficient for the LMM to consistently identify such subtle features, pointing to a need for further specificity or training on similar patterns.
  • 3.
    Prompt 3: The most effective prompt explicitly directs the LMM to examine the four corners of the image for a white square. This approach led to a notable increase in performance metrics, with an accuracy of 77.20% and F1-Score of 71.07%. The improved recall of 56.00%, though better, still indicates a gap in detecting all ’Triggered’ instances. The significant improvement with this prompt underscores the LMM’s capability to focus on specific image areas when instructed. However, the task’s visual simplicity juxtaposed against the performance metrics suggests a limitation in the model’s current ability to parse and analyze image-based details, even with explicit direction. This limitation raises questions about the model’s applicability in more complex or nuanced image-based security tasks without further refinement.
A summary of these results is shown in Figure 5. This analysis of the three prompt types reveals crucial insights into the strengths and limitations of prompt-engineered LMMs in image-based cybersecurity tasks. While tailored prompts can enhance performance, there remains a substantial margin for improvement, especially in tasks requiring detailed visual analysis.
@@ 5.2.1 Prompt engineering with Gemini-pro
Figure 5: Performance of the three prompts in detecting triggers
@@ 5.3 Task 2: Malware Classification
5.3.1 Prompt engineering with Gemini-pro
We use the flow outlined in the methodology section and apply it to the Gemini-pro model. We vary the input prompts to observe how this affects the results. Notably, we investigate three prompt templates to categorize malware images, as shown in Figure 6.
@@ 5.3.1 Prompt engineering with Gemini-pro
  • 1.
    Prompt 1: Starting with a baseline prompt, we ask the LMM to classify images into one of 25 malware types. The model consistently identifies images as belonging to the same class, irrespective of the input. This behavior indicates a significant limitation in the LMM’s ability to distinguish between various malware types. This challenge arises possibly due to the lack of highly specific and contextual information in the prompt, essential for such a fine-grained classification task.
  • 2.
    Prompt 2: Addressing the limitation observed in Prompt 1, we simplify the task by focusing on classifying into five malware families instead of 25 types. The confusion matrix, as depicted in Figure 7, reveals that even with the reduced complexity, the LMM faces difficulties in accurate classification, particularly confusing ’Trojan’ with ’Adware’ and ’Virus’ with ’Backdoor’. This struggle suggests that the task’s inherent complexity and the LMM’s limitations in deriving sufficient context from prompts alone pose significant challenges. The nature of visual data in malware images, which may not be distinctly differentiable through prompts, could be contributing to this challenge.
  • 3.
    Prompt 3: In an attempt to improve classification accuracy, we introduce example images in the prompt for each malware family, asking the model to classify new images based on similarity. The confusion matrix in Figure 8 shows a marginal improvement in performance. However, the results still highlight the inherent difficulties faced by LMMs in processing and classifying complex visual data such as malware images. This suggests that while context and examples aid in performance, they are not sufficient to overcome the LMM’s challenges in handling detailed and nuanced visual classifications.
Figure 7: Confusion matrix for Gemini-pro with prompt 2
Figure 8: Confusion matrix for Gemini-pro with prompt 3
@@ 5.3.2 Fine-tuning ViT to Predict Malware Classes
We fine-tune a ViT model to predict the 25 malware classes of the dataset. We train on the training set, and we report the results on the test set. The model achieved an accuracy of 97.12%, a precision of 97.17%, a recall of 97.11%, and an F1-Score of 97.11%. The high performance of the fine-tuned ViT model in accurately classifying malware classes is indicative of its robust pattern recognition capabilities. The confusion matrix shown in Figure 9 reveals that the model is accurately predicting each of the 25 malware categories. This result is particularly notable, considering the complexity and variety of malware signatures, highlighting the effectiveness of fine-tuning in enhancing model specificity for intricate classification tasks.
@@ 6 Discussion
The results from our experiments offer new insights into the capabilities and limitations of prompt-engineered LMMs and fine-tuned ViTs for cybersecurity applications. This discussion section aims to delve into a comparative analysis of both approaches across the two tasks, underlining the inherent challenges associated with both visually evident and non-visually evident tasks in the context of LMMs and ViTs.
Our experiments in malware classification, a non-visually evident task, highlight a significant challenge for LMMs like Gemini-pro. The limited success in classifying malware types and families, even with enhanced prompts and example images, underscores the difficulty LMMs face in tasks requiring deep contextual understanding and intricate pattern recognition. As illustrated in Figures 7 and 8, the Gemini-pro model struggled to differentiate between various malware classes, indicating a gap in its ability to interpret complex visual data. This suggests that while LMMs can be powerful tools for certain types of image analysis, their effectiveness diminishes in tasks where visual cues are less obvious or require sophisticated interpretation.
Contrary to expectations, the results from the trigger detection task reveal that even visually evident tasks can pose challenges for LMMs. While the Gemini-pro model demonstrated improvement with more detailed prompting, as seen in the increased accuracy from Prompt 1 to Prompt 3, its performance still fell short of perfection, as evidenced by the results reported in Section 6. This indicates that the effectiveness of LMMs in visually evident tasks is highly contingent on the precision and context provided in the prompts. The inability of the model to achieve full accuracy, despite the triggers being visually evident, raises questions about the applicability of LMMs in scenarios where rapid and accurate detection of visual anomalies is critical.
In contrast, the fine-tuned ViTs demonstrated exceptional performance in both tasks. The ViTs achieved perfect scores in the visually evident trigger detection task and remarkable accuracy in the more complex malware classification task, as shown in Figures 9 and 10. These results underscore the fine-tuned ViTs’ robustness in handling both visually evident and non-visually evident tasks, owing to their ability to focus on and analyze specific regions of an image. The success of fine-tuned ViTs in these experiments suggests their better suitability for cybersecurity applications, where the accuracy and reliability of visual analysis are paramount.
@@
In conclusion, our comparative analysis reveals that while LMMs are easier to use and offer a wide range of applicability, they do not perform well in all circumstances. This is particularly evident in tasks requiring high-level visual comprehension or intricate pattern recognition, where LMMs, even with prompt engineering, may fall short. As cybersecurity threats continue to evolve, the adoption of advanced vision models like ViTs, specifically fine-tuned for dedicated tasks, could prove instrumental in developing more robust and intelligent defense mechanisms against a wide array of cyber threats. The superior performance of ViTs in both visually evident and non-visually evident tasks highlights their potential as a reliable and effective solution for diverse cybersecurity challenges.
@@ 7 Conclusion
This study conducted a comparative analysis of prompt-engineered LMMs and fine-tuned ViT models in two distinct cybersecurity tasks: trigger detection (visually evident) and malware classification (non-visually evident). Employing a methodical approach that included prompt engineering for the Gemini-pro model and fine-tuning of pre-trained ViTs, the study aimed to assess the effectiveness of these models in specialized visual tasks.

### v2

@@ 5.1.3 Evaluation Metrics
The performance of LMMs and fine-tuned ViT models was assessed using standard classification metrics, including accuracy, precision, recall, F1-score, and confusion matrix. These metrics provided a comprehensive understanding of each model’s efficacy in correctly identifying triggered images and classifying malware types and families.
@@
  • Recall measures the proportion of actual positives correctly identified.
@@ 5.2 Task 1: Trigger Detection
5.2.1 Prompt engineering with LMMs
We implement the methodology outlined earlier for the five LMMs on the MNIST dataset, experimenting with different input prompts to assess their impact on the models’ performances. Our investigation focuses on three distinct prompt templates for detecting triggers in images, as illustrated in Figure 4.
@@ 5.2.1 Prompt engineering with LMMs
  • Prompt 1: This baseline prompt tasked each LMM with classifying images as ’Triggered’ or ’Unaltered’ without additional details. Here, all models exhibit poor performance as shown in Figure 7.
    • LLaVA showed modest accuracy at 51%, but with a very low recall of 3.2% and an F1-score of 6.1%, indicating severe under-detection of triggered cases despite a precision of 72.7%.
    • BakLLaVa had a similar accuracy at 51.6%, but exhibited a much higher recall of 66.4%, balanced by a lower precision of 51.2%, resulting in an F1-score of 57.8%.
    • Moondream 2 classified all images as triggered, achieving 50% accuracy and precision, with a perfect recall of 100%, resulting in an F1-score of 66.7%.
    • Gemini-pro-vision predicted all images as unaltered, which resulted in a 50% accuracy and 0% recall. The precision and F1-score could not be calculated here, because no instances were classified as triggered.
    • GPT-4o had a low accuracy of 52.4%, with extremely high recall of 98.3% and a high precision, resulting in an F1-score of 65.9%.
  • Prompt 2: This prompt included a specific description of the trigger as a small white square to guide the LMM on what to look for in the input image. As a result, the models that performed better are Gemini-pro-vision and GPT-4o, as shown in Figure 7:
    • LLaVA’s accuracy slightly decreased to 49.4%, with low recall and precision, and an F1-score of 10.6%.
    • BakLLaVa showed a further decrease in accuracy to 46.2%, with improved precision to 42% and an F1-score of 27%.
    • Moondream 2 saw a slight accuracy increase to 51%, with high recall (79.6%) but moderate precision (50.6%), resulting in an F1-score of 61.9%.
    • Gemini-pro-vision achieved a notable accuracy increase to 63%, with precision jumping to 87.4%, but a lower recall of 30.4% suggested missed ’Triggered’ cases, leading to an F1-score of 45.1%.
    • GPT-4o excelled under this prompt with an accuracy of 91%, high precision (98.9%), and substantial recall (81.6%), resulting in an F1-score of 89.4%.
  • Prompt 3: This prompt explicitly directed the models to examine the four corners of the image for a white square in an effort to direct the attention of the LMM to specific parts of the image. The results of applying this prompt to each LMM are shown in Figure 7:
    • LLaVA 1.5’s performance remained low with an accuracy of 50.2%, and extremely low recall of 2.4%, resulting in an F1-score of 4.6%.
    • BakLLaVa’s accuracy was slightly better at 47.4%, with a moderate increase in recall to 27.6% and an F1-score of 34.4%.
    • Moondream 2 showed marginal improvement with accuracy at 52.6% and high recall at 75.6%, leading to an F1-score of 61.5%.
    • Gemini-pro-vision’s accuracy soared to 77.2%, with high precision (97.2%) and more than half the cases correctly recalled (56%), resulting in an F1-score of 71.1%.
    • GPT-4o again performed the best with an accuracy of 91.9%, very high precision (95.3%), and an 87.1% recall rate, leading to an F1-score of 91%.
This detailed analysis reveals crucial insights into the strengths and limitations of prompt-engineered LMMs in image-based cybersecurity tasks. The results show that while tailored prompts can enhance performance significantly for some LMMs like Gemini-pro-vision and GPT-4o, other models still encounter challenges and show no clear improvement.
Figure 5: Performance of each LMM with prompt 1
Figure 6: Performance of each LMM with prompt 2
Figure 7: Performance of each LMM with prompt 3
@@ 5.3 Task 2: Malware Classification
5.3.1 Prompt engineering with LMMs
Since in the visually-evident task, only Gemini-pro-vision and GPT-4o showed acceptable performance, these are the only models that will be used for the malware classification task since it is more complex. We use the flow outlined in the methodology section and apply it to the selected LMMs. We vary the input prompts to observe how this affects the results. Notably, we investigate three prompt templates to categorize malware images, as shown in Figure 8.
@@ 5.3.1 Prompt engineering with LMMs
  • Prompt 1: Starting with a baseline prompt, we ask the LMM to classify images into one of 25 malware types. The two models consistently identifiy images as belonging to the same class, irrespective of the input. Gemini-pro-vision always predicted "Fasong" as the malware class, and GPT-4o always predicted "Allaple" as the malware class. This behavior indicates a significant limitation in the LMM’s ability to distinguish between various malware types. This challenge arises possibly due to the lack of highly specific and contextual information in the prompt, essential for such a fine-grained classification task.
  • Prompt 2: To address the limitations observed in Prompt 1, we simplified the task by focusing on classifying into five malware families instead of 25 types. Despite this reduction in complexity, the models still exhibit a bias towards specific categories, although they occasionally predict other categories as well. The confusion matrices, as depicted in Figure 9, show that even with simplified criteria, the LMMs struggle with accurate classification. Specifically, Gemini-pro-vision shows a pronounced bias towards the "Adware" class, while GPT-4o tends towards the "Trojan" class. The overall accuracies of Gemini-pro-vision and GPT-4o are 21.2% and 20.4%, respectively, with their macro F1-scores being 11.7% and 7.61%, respectively. This difficulty suggests that the task’s inherent complexity and the LMMs’ limitations in extracting sufficient context from prompts alone present significant challenges. Additionally, the nature of visual data in malware images, which may not be distinctly differentiable through prompts alone, could be contributing to these challenges.
  • Prompt 3: In an attempt to improve classification accuracy, we introduce example images in the prompt for each malware family, asking the model to classify new images based on similarity. The confusion matrix in Figure 10 shows a marginal improvement in performance. However, there is still some bias towards a specific class: "Worm" in the case of Gemini-pro-vision and "Trojan" in the case of GPT-4o. Gemini-pro-vision achieves an accuracy of 17.6% and a macro F1-score of 12.7%, whereas GPT-4o achieves an accuracy of 33.6% and a macro F1-score of 34.63%. Even if GPT-4o shows more improvement compared to Gemini-pro-vision, the results still highlight the inherent difficulties faced by LMMs in processing and classifying complex visual data such as malware images. This suggests that while context and examples aid in performance, they are not sufficient to overcome the LMM’s challenges in handling detailed and nuanced visual classifications.
Figure 9: Confusion matrices for Gemini-pro-vision and GPT-4o with prompt 2
Figure 10: Confusion matrices for Gemini-pro-vision and GPT-4o with prompt 3
@@ 5.3.2 Fine-tuning ViT to Predict Malware Classes
We fine-tune a ViT model to predict the 25 malware classes of the dataset. We report the results on the same test set we used to evaluate LMMs. The model achieved an accuracy of 97.12%, a precision of 97.17%, a recall of 97.11%, and an F1-Score of 97.11%. The high performance of the fine-tuned ViT model in accurately classifying malware classes is indicative of its robust pattern recognition capabilities. The confusion matrix shown in Figure 11 reveals that the model is accurately predicting each of the 25 malware categories. This result is particularly notable, considering the complexity and variety of malware signatures, highlighting the effectiveness of fine-tuning in enhancing model specificity for intricate classification tasks.
@@ 6 Discussion
The results from our experiments offer new insights into the capabilities and limitations of prompt-engineered LMMs and fine-tuned ViTs for cybersecurity applications. In this section, we perform a comparative analysis of both approaches across the two tasks, underlining the inherent challenges associated with both visually evident and visually non-evident tasks in the context of LMMs and ViTs.
The results from the trigger detection task reveal that even visually evident tasks can pose challenges for LMMs. While models like Gemini-pro-vision and GPT-4o demonstrated improvement with more detailed prompting, as seen in the increased accuracy from Prompt 1 to Prompt 3, their performance still fell short of perfection, as evidenced by the results reported in Section 5. This indicates that the effectiveness of LMMs in visually evident tasks is highly contingent on the precision and context provided in the prompts. The inability of the model to achieve full accuracy, despite the triggers being visually evident, raises questions about the applicability of LMMs in scenarios where rapid and accurate detection of visual anomalies is critical. Moreover, through the experiments, it was demonstrated that even with prompt engineering, not all LMMs exhibit improvement. For instance, LLaVA, BakLLaVA, and moondream2 show almost a constant performance across the three prompts used.
On the other hand, our experiments in malware classification, a visually non-evident task, highlight a significant challenge for LMMs. The limited success in classifying malware types and families, even with enhanced prompts and example images, underscores the difficulty LMMs face in tasks requiring deep contextual understanding and intricate pattern recognition. As illustrated in Figures 9 and 10, while prompt engineering offers some improvement, the models struggled to differentiate between various malware classes, indicating a gap in their ability to interpret complex visual data. This suggests that while LMMs can be powerful tools for certain types of image analysis, their effectiveness diminishes in tasks where visual cues are less obvious or require sophisticated interpretation.
In contrast, the fine-tuned ViTs demonstrated exceptional performance in both tasks. The ViTs achieved perfect scores in the visually evident trigger detection task and remarkable accuracy in the more complex malware classification task, as shown in Figures 11 and 12. These results underscore the fine-tuned ViTs’ robustness in handling both visually evident and non-evident tasks, owing to their ability to focus on and analyze specific regions of an image. The success of fine-tuned ViTs in these experiments suggests their better suitability for cybersecurity applications, where the accuracy and reliability of visual analysis are paramount.
@@
In conclusion, our comparative analysis reveals that while LMMs are easier to use and offer a wide range of applicability, they do not perform well in all circumstances. This is particularly evident in tasks requiring high-level visual comprehension or intricate pattern recognition, where LMMs, even with prompt engineering, may fall short. As cybersecurity threats continue to evolve, the adoption of advanced vision models like ViTs, specifically fine-tuned for dedicated tasks, could prove instrumental in developing more robust and intelligent defense mechanisms against a wide array of cyber threats. The superior performance of ViTs in both visually evident and visually non-evident tasks highlights their potential as a reliable and effective solution for diverse cybersecurity challenges.
@@ 7 Conclusion
This study conducted a comparative analysis of prompt-engineered LMMs and fine-tuned ViT models in two distinct cybersecurity tasks: trigger detection (visually evident) and malware classification (visually non-evident). Employing a methodical approach that included prompt engineering for five LMMs and fine-tuning of pre-trained ViTs, the study aimed to assess the effectiveness of these models in specialized visual tasks.
@@
Acknowledgements
The authors would like to acknowledge that this work has been supported by the Maroun Semaan Faculty of Engineering and Architecture (MSFEA) at the American University of Beirut (AUB)